OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <8, 1, -6>; v = <6, 7, -3>; w = <-7, 5, 2>

OpenStudy (anonymous):

dot product, multiply element by element and add... 6(-7)+7(5)+(-3)(2)=-42+35-6=-13 that is if my calculations are right

OpenStudy (anonymous):

and that would just be the final answer? simple as that? @josiahh

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thanks. are you 100% sure its -13?

OpenStudy (anonymous):

well check my calculations if u want

OpenStudy (anonymous):

just did, your right! :) thanks!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Find a ⋅ b. a = <2, 4>, b = <2, 5>

OpenStudy (anonymous):

<4, 9> 24 <4, 20> 16

OpenStudy (anonymous):

@josiahh

OpenStudy (anonymous):

do ur own work

OpenStudy (anonymous):

btw, feather its 4,20 like weed